Dolphins varied directly as tuna and inversely as sharks squared. When there were 50 dolphins, there were 125 tuna and 5 sharks. How many dolphins were there when there were 16 tuna and only 2 sharks?
Answer by greenestamps(13200)



The number of dolphins varies directly as the number of tuna. The number of tuna changes by a factor of 16/125, so the number of dolphins changes be a factor of 16/125.

The number of dolphins varies inversely as the square of the number of sharks. The number of sharks changes by a factor of 2/5, so the number of dolphins changes by a factor of (5/2)^2 = 25/4.

ANSWER: (50)(16/125)(25/4) = 40